# Linying Lv

### v3

@@ Abstract
I examine the value of information from sell-side analysts by analyzing a large corpus of their written reports. Using embeddings from state-of-the-art large language models, I show that textual information in analyst reports explains 10.19% of contemporaneous stock returns out-of-sample, a value that is economically more significant than quantitative forecasts. I then perform a Shapley value decomposition to assess how much each topic within the reports contributes to explaining stock returns. The results show that analysts’ income statement analyses account for more than half of the reports’ explanatory power. Expressing these findings in economic terms, I estimate that early acquisition of analysts’ reports can yield significant profits. Analysts’ information value peeks in the first week following earnings announcements, highlighting their vital role in interpreting new financial data. JEL Classification: G11, G14, G24
@@ 1 Introduction
Sell-side analysts play a crucial role in financial markets by producing and processing information for investors. Their research reports, which include both quantitative forecasts and qualitative analyses, are widely used by investors in decision-making (Barber et al., 2001; Kong et al., 2021). This raises a fundamental question: Do the analysts generate value for their clients, or are they merely peddling expensive noise?
@@
Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of textual data. Prior studies have shed light on the incremental information value of report text through sentiment analyses (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018). However, analyst reports are meant to provide contextual information beyond the simple sentiment dimension. The context has been perceived at least as important as the quantifiable outputs. According to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations. Yet, empirical evidence supporting this survey feedback is limited, and it remains unclear what specific discussion holds the most value for investors.
@@
The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning beyond sentiment and topics, allowing us to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2024). Consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences analyze earnings in a negative tone, the latter provides additional contextual information and reasoning. The question now becomes: do investors disregard obvious facts and figures in favor of details in lines?
@@
I investigate the question through three general steps: analyst output representation, econometric modeling, and value decomposition. Analyst outputs include both quantitative information and qualitative information. The quantitative data consists of numerical forecasts like stock recommendations, earnings forecasts, and target prices. For qualitative textual information, I extract contextualized representations by mapping each report into a structured LLM embedding space.
@@
With structured representation, I evaluate how well analyst information explains stock returns as evidence of information content. A significant challenge in projecting stock returns onto text embeddings is the risk of in-sample overfitting due to high dimensionality. To address this issue, I employ Ridge regressions. The simple model penalizes large coefficients while still preserving the model’s ability to capture essential signals that move the market. Additionally, I use out-of-sample R-squared to evaluate model performances.
@@
I find that the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR_{[-1,+1]}$), surpassing the 9.01% explained by numerical information. When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using either information type alone. These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the distinct and economically meaningful information content embedded in analyst report narratives.
@@
The informativeness of analyst reports varies substantially with forecast changes. Analyst reports accompanied by forecast revisions provide significantly more valuable information to financial markets compared to reiterations. For reports containing recommendation revisions, the combination of numerical and textual information generates an out-of-sample $R^{2}$ of 22.63%. In contrast, reports merely reiterating previous earnings forecasts yield statistically insignificant out-of-sample $R^{2}$ is not statistically significant, indicating they provide minimal new information to market participants.
@@
Previous research emphasizes analysts’ role in information processing around corporate earnings announcements( e.g., Livnat and Zhang, 2012, Keskek et al., 2014, Kim and Song, 2015, Lobo et al., 2017, and Barron et al., 2017). My analysis on samples during and outside earnings announcement periods reveals that analyst reports’ information content peaks in the first week following earnings releases. The finding aligns with Huang et al. (2018), who argue that prompt responses to corporate disclosures provide valuable insights to clients. Using distinct models trained on samples during and outside earnings announcement windows, the out-of-sample $R^{2}$ approximately doubles in the first week following earnings announcement dates compared to other periods. The substantial increase highlights the incremental value analysts add by interpreting and contextualizing earnings information for investors.
@@
To address the concern that my results merely capture analysts’ “piggybacking” on earnings conference call information or post-earning-announcement drift (PEAD), I control for both earnings surprises and the latest earnings conference call transcripts using text embeddings. The results show that analyst reports provide distinct and valuable insights even after accounting for these alternative explanations.
As raised in the review paper of Bradshaw et al. (2017), identifying the qualitative attributes of analyst reports that investors find most salient remains largely unsolved. I develop a systematic approach to quantify the importance of different report content. By exploiting the additive property of text embeddings and implementing a Shapley value decomposition, I attribute the explanatory power to 17 major topics discussed in analyst reports.
@@
To systematically quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports. The results reveal that Income Statement Analyses, particularly interpretations of realized earnings, account for over half of the reports’ explanatory power in terms of out-of-sample $R^{2}$. These results reinforce the critical role of analysts in interpreting financial data and constructing the narratives and context essential for investment decisions.
@@
To translate statistical measures of information content into economic terms, I adopt Kadan and Manela (2024)’s framework to quantify the dollar value of analyst information for strategic investors. The multi-client framework provides a realistic foundation for analyzing the complex dynamics of analyst reports dissemination. The information value can be interpreted as the total expected profits for strategic investors either because: (1) some investors have early access to analyst views, or (2) they conduct independent research and derive insights comparable to those of analysts. Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to price impact, representing the cost of trading on this information.
@@
The analysis reveals substantial economic value in analyst information. For an average S&P 100 stock over a three-day window, strategic investors’ expected profits from early access to analyst reports amount to \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million when combined. Based on an average of 15 report days annually, the aggregate annualized information value reaches \$6.89 million—a conservative lower bound according to Kadan and Manela (2024). This value increases for large-cap stocks, bold analysts, and reports released promptly following earnings announcements.
My paper relates to three strands of literature. It first connects to recent studies of market reactions to security analyst reports (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024). Earlier investigations primarily rely on manual text coding and sentiment analysis, which are labor-intensive and may overlook nuanced insights. My work demonstrates the superiority of state-of-the-art LLMs in generating contextualized textual representations. Compared to Huang et al. (2014)’s Naive Bayes tone classification, LLMs offer two main advantages: better capture of overall report sentiment and extraction of richer information beyond tone. Moving from simple tone analysis to comprehensive LLM embeddings boosts the $R^{2}$ from 3% to over 10%. The substantial improvement in explanatory power highlights LLMs’ potential to transform textual analysis. Additionally, the topic decomposition approach provides novel insights into analysts’ value by revealing their role in interpreting financial data.
Second, this work contributes to the growing field of machine learning and artificial intelligence in finance. Recent studies like Chen et al. (2022), Li et al. (2024), Jha et al. (2024), and Beckmann et al. (2024) have showcased the potential of implementing LLMs in a financial paradigm. However, the interpretability of LLMs remains an issue, and solutions like prompting can introduce subjectivity and manipulation concerns. By exploiting the additive property of LLM embeddings and developing a Shapley value decomposition approach, my research provides a new framework for explainable AI.
Third, this paper advances the understanding of analyst tipping - the practice of brokerage firms providing early research access to high-commission clients (Irvine et al., 2007; Christophe et al., 2010). Tipping allows brokerages to offset research costs by earning commissions from clients who benefit from the privileged information. Prior studies, such as those by Green (2006) and Kadan et al. (2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%. This paper extends by quantifying the economic dollar value of such privileged information.
@@
The rest of the paper is organized as follows. Section 2 introduces methodology, measures of analyst information value, and interpretation framework. Section 3 describes the data and empirical results. Section 4 concludes.
@@ 2 Methodology
This section describes the key methodologies: the use of large language models for topic modeling and text embeddings, estimation of information content and strategic information value, and decomposition metrics for evaluating topic-specific information value. [^1] The details of machine learning models used to explain contemporaneous stock returns are provided in the Online Appendix C.
2.1 Large Language Models
The analysis of analyst reports leverages three advanced generative AI models. A fine-tuned BERT model classifies sentences according to predefined criteria, ChatGPT extracts and synthesizes major topics through prompting, and LLaMA generates high-dimensional vector embeddings to represent report content.
2.1.1 ChatGPT Prompting
To avoid ad-hoc classification, I employ ChatGPT prompting to extract exclusive and meaningful topics. The goal is to identify primary topics covered in analyst reports and assign each sentence to a single topic category.
@@ 2.1.1 ChatGPT Prompting
In the first step, I use a random sample of 100 analyst reports and request ChatGPT to analyze the content and identify high-level categories with the following prompt:
@@
Prompt 1:Please read the provided text file of sell-side analyst reports carefully. What are high-level mutually exclusive topics covered in these reports? Make sure that each sentence from the text file can be assigned to one of the topics. Here is the report content: {text}.
@@
It ends up with 16 categories defined in Table LABEL:tab:_topic. A category labeled “None of the Above” is included to accommodate sentences that do not fit into predefined topics. To provide a visual representation of the content within each topic, Figure A1 in the Online Appendix presents word clouds for each category. Furthermore, Table LABEL:tab:_topic_example offers illustrative sentences extracted directly from analyst reports, providing concrete examples of the discussions typical to each category.
@@
In the second step, a systematic approach is required to assign topic classifications to a large corpus of 6,975,114 sentences. To begin with, I employ the GPT-4-Turbo model to categorize each sentence in a random sample of 17,028 sentences from 100 analyst reports into a single, most relevant topic.
@@
This labeled sample is used to fine-tune a BERT model, which then classifies all sentences in the corpus. The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample. The distribution of topics is relatively stable across years, as illustrated in Figure 1. Income Statement Analyses and Financial Ratios are the most frequently discussed topics in analyst reports, accounting for 17.23% and 15.65% of all sentences, respectively. These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics. Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least frequent topics.
Several factors contribute to these observed frequencies. The low occurrence of Appendices and Disclosures can be attributed to pre-boilerplate cleaning procedures implemented in the data preparation phase. Executive Summary sections appear infrequently because they are difficult to identify at the sentence level. While ESG discussions are minimal overall, they increase notably after 2020, reflecting growing market interest in sustainability issues.
2.1.2 Text Representations
I need a structured representation of writtrn reports. Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps. Prior to tokenization, the text undergoes normalization, which may include converting to lowercase, removing punctuation, and handling special characters. Depending on the strategy, tokenization involves breaking down text into smaller units, which can be words, subwords, or even characters. These tokens are then mapped to unique identifiers based on a predefined vocabulary learned during model training.
To illustrate this process, consider the following sentence from an analyst report on NVIDIA: “We expect to see more GPU/SmartNIC integration as next-gen workloads grow and CPUs become a bigger bottleneck in the data center.” The LLaMA tokenizer breaks this into 34 sub-word tokens, including “‘We’, ‘expect’, ‘to’, ‘see’, ‘more’, ‘GPU’, ‘/’, ‘Sm’, ‘art’, ‘N’, ‘IC’, ‘integration’, ‘as’, ‘next’, ‘-’, ‘gen’, ‘work’, ‘loads’, ‘grow’, ‘and’, ‘CPU’, ‘s’, ‘become’, ‘a’, ‘bigger’, ‘bott’, ‘l’, ‘ene’, ‘ck’, ‘in’, ‘the’, ‘data’, ‘center’, ‘.’”. This granular approach allows for nuanced interpretation of complex terms. For instance, ‘next-gen workloads’ becomes five separate tokens: ‘next’, ‘-’, ‘gen’, ‘work’, and ‘loads’.
At the heart of LLMs lies the transformer architecture. A core output of the transformer is text embeddings - dense, continuous vectors that capture semantic and syntactic information. Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia. These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
For this analysis, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI. The input length limit of LLaMA-2 models is 4,096 tokens, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens. The resulting contextual embeddings capture nuanced report information and are suitable for downstream tasks like sentiment analysis and economic modeling.
2.2 Information Content

### v4

@@ Abstract
I examine the value of information from sell-side analysts by analyzing a large corpus of their written reports. Using embeddings from state-of-the-art large language models, I show that qualitative information in analyst reports explains above 10% of contemporaneous stock returns out-of-sample, a value that is more economically significant than quantitative forecasts. I then perform a Shapley value decomposition to assess how much each topic within the reports contributes to explaining stock returns. The results show that analysts’ income statement analyses account for more than half of the reports’ explanatory power. Expressing these findings in economic terms, I estimate that early acquisition of analyst reports can yield significant profits. Analyst information value peaks in the first week following earnings announcements, highlighting their vital role in interpreting new financial data. JEL Classification: G11, G14, G24
@@ 1 Introduction
Sell-side analysts play a crucial role in financial markets by producing and processing information for investors. Their research reports, which include both quantitative forecasts and qualitative analyses, are widely used by investors in decision-making (Barber et al., 2001; Kong et al., 2021). This raises a fundamental question: Do analysts generate value for their clients, or are they merely peddling expensive noise?
@@
Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of text. Prior studies have shed light on the incremental information value of report text through sentiment analysis (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018). However, analyst reports are meant to provide contextual information beyond the simple sentiment dimension. The context has been perceived to be at least as important as the quantitative outputs. According to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations. Yet, empirical evidence supporting this survey feedback is limited, and it remains unclear what specific discussion holds the most value for investors.
@@
The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning beyond sentiment and topics, allowing us to capture both contextual information and the reasoning logic within the text (see, for example: Chen et al., 2022; Li et al., 2024). Consider the following two statements: “Reported EPS of \$0.94 misses consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences analyze earnings in a negative tone, the latter provides additional contextual information and reasoning. The question now becomes: When presented with both, do investors prioritize the explicit figures or the explanatory narrative?
@@
I investigate the research question through three general steps: analyst output representation, econometric modeling, and value decomposition. For quantitative information, I consider three numerical forecasts, including stock recommendations, earnings forecasts, and target prices. For qualitative information, I extract contextualized representations by mapping each report into a structured LLM embedding space.
@@
With structured representation, I evaluate how well analyst information explains stock returns as evidence of information content. Due to the high dimensionality of LLM embeddings, projecting stock returns on them bears overfitting risk. To address this issue, I employ the Ridge regression. The simple linear model penalizes large coefficients while still preserving the model’s ability to capture essential signals that move the market.
@@
I find that the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR_{[-1,+1]}$), surpassing the 9.01% explained by numerical information. When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using either information type alone. These findings remain robust across various LLM representations, machine learning algorithms, and $CAR$ windows, highlighting the distinct and economically meaningful information content embedded in analyst report narratives.
@@
The informativeness of analyst reports varies substantially depending on whether they are accompanied by forecast revisions. Reports that revise earnings forecasts or recommendations convey significantly more value to investors than those that simply reiterate prior views. For reports containing recommendation revisions, the combination of numerical and textual information generates an out-of-sample $R^{2}$ of 22.63%. In contrast, reports merely reiterating previous earnings forecasts yield statistically insignificant out-of-sample $R^{2}$.
@@
Previous research emphasizes analysts’ role in processing information around corporate earnings announcements (e.g., Livnat and Zhang, 2012, Keskek et al., 2014, Kim and Song, 2015, Lobo et al., 2017, and Barron et al., 2017). Consistent with this literature, my analysis shows that the information content of analyst reports peaks in the first week following earnings releases. Comparing models trained separately on post-earnings and non-earnings periods, I find that the out-of-sample $R^{2}$ approximately doubles during the one-week earnings announcement window. The substantial increase highlights the incremental value analysts add by interpreting and contextualizing earnings news for investors.
@@
To address the concern that the explanatory power may simply reflect “piggybacking” on earnings conference call information or post-earnings-announcement drift (PEAD), I control for both earnings surprises and the latest conference call transcripts. The results confirm that analyst reports provide distinct and valuable insights.
@@
An unresolved question in the literature, highlighted by Bradshaw et al. (2017), is which qualitative aspects of analyst reports investors find most informative. This study addresses this gap by developing a systematic approach to quantify the relative importance of various types of report content. Specifically, I leverage the additive structure of text embeddings and implement Shapley value decomposition to attribute the explanatory power of analyst reports across 17 prevalent topics. The results reveal that Income Statement Analyses, especially interpretations of realized earnings, account for over half of this explanatory power. These results reinforce analysts’ critical role in interpreting financial data and constructing the narratives essential for investment decisions.
@@
To translate the statistical informativeness of analyst reports into economic significance, I adopt the Kadan and Manela (2025) framework to quantify the dollar value of analyst information for strategic investors. Their multi-agent model provides a realistic basis for assessing the economic impact of analyst report dissemination. The information value represents the total expected profits accruing to strategic investors who either: (1) gain early access to analyst views, or (2) independently derive insights comparable to those from analysts. Intuitively, the measure reflects both explainable return variance (the uncertainty reduction from analyst information) and the price impact (the cost of trading on such information).
@@
The analysis indicates substantial economic value in analyst information. For an average S&P 100 stock, the estimated profit over a three-day window for strategic investors with early access (two days prior to public release) is approximately \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million when both sources are combined. Assuming an average of 15 report days per year, this implies an annualized information value of \$6.89 million, a conservative lower bound according to Kadan and Manela (2025). The value is notably higher for large-cap stocks, reports from bold analysts, and those released promptly following earnings announcements.
My paper relates to three strands of literature. First, this paper offers several extensions to prior research that highlight the incremental information in report narratives and the interpretative role of analysts (e.g., Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018). Specifically, I show that more contextualized textual representations not only provide significant incremental information but can also surpass the explanatory power of analysts’ traditional quantitative forecasts. Additionally, I demonstrate that the informational value derived from the report text extends substantially beyond simple sentiment measures. To illustrate, transitioning from basic textual tone measures to comprehensive LLM embeddings increases the out-of-sample $R^{2}$ from approximately 3% to over 10%.
Furthermore, to my knowledge, this is the first study to systematically identify and quantify the specific qualitative attributes in analyst reports that investors find most salient. The research complements the existing literature by providing a more granular understanding of which textual elements drive the value of analyst reports and when that value peaks, such as after earnings announcements or when accompanying forecast revisions.
Second, this paper makes a distinct methodological contribution to the expanding literature on AI and machine learning in finance. While the application of LLMs in financial paradigms is promising (e.g., Chen et al., 2022; Li et al., 2024; Lopez-Lira and Tang, 2023; Jha et al., 2024), their inherent ‘black-box’ nature presents significant hurdles for model validation and reliability. To address this critical interpretability gap, this paper implements a Shapley value decomposition framework specifically designed for LLM embeddings plus topic modeling. By leveraging the additive structure of these embeddings, this approach systematically attributes an LLM’s performance to constituent textual features, thereby enhancing model transparency and offering a robust pathway toward explainable AI in financial applications.
Third, this paper offers new insights into analyst tipping. It provides a rigorous quantification of the economic dollar value accruing from privileged early access to research, a practice of brokerage firms providing early research access to high-commission clients (Irvine et al., 2007; Christophe et al., 2010). While prior work (e.g., Green, 2006; Kadan et al., 2018) documents significant abnormal returns for clients with early access, this study further translates the information advantages into tangible monetary terms. The dollar value offers a more concrete measure for understanding the incentive of selective disclosure and the scale of wealth transfer involved.
@@
The rest of the paper is organized as follows. Section 2 introduces the methodology, measures of analyst information value, and the interpretation framework. Section 3 describes the data and empirical results. Section 4 concludes.
@@ 2 Methodology
This section describes the key methodologies: using large language models for text representation and topic modeling, and the decomposition techniques to evaluate the information value of individual topics.
2.1 Text Representation
A core step in analyzing the information value of analyst reports is constructing structured, numerical representations of textual content. LLMs offer a significant advancement by providing richer, more contextualized linguistic representations than traditional methods.
As illustrated in Figure A1 in the online Appendix, the process begins with text tokenization, where text segments are mapped to unique identifiers based on a predefined vocabulary. These tokens are then processed through a transformer architecture. The transformer’s key output consists of dense, continuous vectors that capture both semantic and syntactic information. To generate report-level representations, I compute embeddings as the average across all tokens and transformer layers:
$y^{emb}=\frac{1}{N_{k}N_{l}}\sum_{l=1}^{N_{l}}\sum_{i=1}^{N_{k}}{e}_{i,l},$ (1)
where $e_{i,l}$ denotes the contextual embedding of token $i$ generated by layer $l$, $N_{k}$ is the total number of tokens in the report, $N_{l}$ is the number of layers from which embeddings are aggregated.
The embedding representation provides two salient advantages over traditional sentiment measures. First, LLMs position each token within a high-dimensional semantic space. This maintains complex relationships both within and across sentences, producing tone measurements that are substantially more nuanced and reliable than those generated by dictionary-based approaches. Second, these embeddings encode context-dependent meaning and narrative structure, thereby conveying information that extends well beyond simple positive-negative classifications and conveys far richer linguistic signals.
For this study, I use MetaAI’s pre-trained LLaMA-2-13B model and its associated tokenizer. With an input length limit of 4,096 tokens, the model accommodates almost all analyst reports, which have a median length of 1,393 tokens and a mean of 2,055 tokens. The resulting contextual embeddings serve as the foundation for downstream tasks such as topic attribution and explaining stock returns.
2.2 Topic Modeling
This section introduces the methodology used to identify prevalent themes in analyst reports and to classify each sentence into a single, primary topic. The initial challenge is selecting a topic modeling approach that effectively avoids ad-hoc classifications and ensures economically meaningful interpretations. I exclude manual topic construction to minimize researcher subjectivity.
Given the availability of sentence-level embeddings, clustering these embeddings based on semantic proximity is a straightforward procedure to consider. Accordingly, I expriment a standard K-Means algorithm for initial assessment. [^1] Topic classification results from the K-Means experimentation are available upon request. The unsupervised clustering approach is not adopted as the final methodology for two primary reasons. First, due to the industry-sensitive nature of analyst reports, K-Means tended to generate topics heavily aligned with particular sectors (e.g., pharmaceuticals, banking), whereas this study seeks to identify common thematic categories that span across industries. Second, as also noted by Aleti and Bollerslev (2025), topics derived from purely data-driven clustering tend to be messy and difficult to interpret from an economically meaningful perspective.
To provide a set of distinct and interpretable topics, this study uses targeted prompting of the ChatGPT-4o model. The objective of this approach is to identify a comprehensive set of primary topics consistently discussed in analyst reports and to assign each sentence to one of these predefined categories.
@@ 2.2 Topic Modeling
In the first step, I use a random sample of 100 analyst reports and request ChatGPT-4o model to analyze the content and identify high-level categories with the following prompt:
@@
Prompt 1: Please read the provided text file of sell-side analyst reports carefully. What are high-level, mutually exclusive topics covered in these reports? Make sure that each sentence from the text file can be assigned to one of the topics. Here is the report content: {text}.
@@
It ends up with 16 categories defined in Table LABEL:tab:_topic. A category labeled “None of the Above” is included to accommodate sentences that do not fit into any of the 16 topics. Table LABEL:tab:_topic_example offers illustrative sentences of each category.
@@
In the second step, I build a sentence classifier to assign each sentence to one topic. To begin with, I employ the GPT-4-turbo model to categorize each sentence in a random sample of 100 analyst reports into a single, most relevant topic.
@@
I then fine-tune a BERT-based classifier using the labeled sample and classify all 6,975,114 sentences in the sample. The distribution of topics is relatively stable across years, as illustrated in Figure 1. Income Statement Analyses and Financial Ratios are the most frequently discussed topics, accounting for 17.23% and 15.65% of all sentences, respectively. These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics. While ESG discussions are minimal overall (0.23%), they increase notably after 2020, reflecting growing market interest in sustainability issues.
I validate the topic modeling procedure based on three key evidence. First, the fine-tuned BERT classifier demonstrates high performance, achieving 89% accuracy on the test sample. Given the challenge of distinguishing among 17 topic categories, it demonstrates a notably strong performance. Second, the 16 GPT-generated topics provide comprehensive coverage of the report content. Less than 5% of sentences fall into the fallback category “None of the Above”, indicating that the topic set captures nearly all relevant content. Third, Figure A2 presents word clouds showing the most frequent terms within each topic. The word frequencies align closely with expected topic content. For example, “Income Statement Analysis” is strongly associated with terms like revenue, EPS, and sales, confirming the semantic coherence of the classification.
2.3 Shapley Value Decomposition
Next, I propose a methodology to decompose the value of different topics within analyst reports. As a starting point, a simple approach is aggregating the embeddings of tokens corresponding to each topic and measure their corresponding value independently. However, this method has two significant drawbacks. First, the self-attention architecture of transformer models creates cross-topic contamination. It means, each token’s embedding incorporates contextual information from all other tokens in the report, including those assigned to different topics. Second, topics are not independent. Their interactions can be a primary driver of value. For example, the combination of a valuation discussion and an earnings analysis may be more informative than either alone.
To mitigate the first concern of cross-topic contamination, I construct sentence-segmented embeddings. Specifically, each sentence is processed by the LLM individually. The procedure ensures that the resulting token embeddings $\left(e_{i}\right)$ are contextualized only by the words within their own sentence. A report-level embedding $\left(y^{emb}\right)$ is then constructed as the token-weighted average of these sentence embeddings:
$y^{emb}=\sum_{i=1}^{n}\frac{Token_{i}}{\sum_{i=1}^{n}Token_{i}}S_{i}^{emb},$ (2)
where $S_{i}^{emb}$ is the average token embeddings of sentence $i$, $n$ is the number of sentences in the report, and $Token_{i}$ is the number of tokens in sentence $i$.
To account for topic interdependency, I implement the Shapley Additive exPlanations (SHAP) framework. Originating in cooperative game theory (Shapley, 1953), the Shapley value fairly allocates a collective payoff among contributing members. This concept is widely used in machine learning to ensure fair attribution of model predictions to individual features (Lundberg and Lee, 2017; Chen et al., 2022).
The SHAP framework is ideal for topic value attribution for two reasons. First, it satisfies the additivity (or efficiency) property, ensuring that the total information content of a report is fully distributed among all constituent topics without omissions or double-counting. Second, it inherently accounts for topic interdependencies by evaluating the marginal contribution of each topic across all possible topic combinations.
The Shapley value $\left(\varphi_{p}\right)$ for a specific topic $p$, which quantifies its contribution to the overall out-of-sample explanatory power ($R_{\text{oos }}^{2}$), is defined as:
$\varphi_{p}(R_{oos}^{2})=\sum_{S\subseteq P\setminus{p}}\frac{|S|!(|P|-|S|-1)!}{|P|!}\left[R_{oos}^{2}(S\cup{p})-R_{oos}^{2}(S)\right],$ (3)
where $P$ is the set of all topics in a report. The summation is over all subsets $S$ of topics excluding $p$. The term $R_{\text{oos }}^{2}(S)$ denotes the out-of-sample $R^{2}$ of a model whose input is the report embedding constructed using only sentences belonging to topics in the subset $S$ [^2] I define missing topic embeddings as zero vectors, effectively removing them from the model input. For a given subset of topics $S$, the corresponding embedding is constructed using the token-weighted average formula above, but summed only over sentences classified into topics within $S$.
3 Data and Empirical Results
This section describes the data and presents empirical results of the analyst information value. I begin with an examination of analyst information content and its decomposition across topics. I then translate these findings into economic terms by quantifying the strategic dollar value of analyst information.
3.1 Data
The initial sample comprises 223,091 sell-side analyst reports of S&P 100 constituent stocks from 2000 to 2023 downloaded from the Mergent Investext database. After converting the reports from their original PDF format to plain text using Adobe Acrobat, the extracted text is segmented into individual sentences. To focus on stock-relevant discussions, I fine-tune a BERT classifier to identify and remove boilerplate content. [^3] Boilerplate segments typically contain disclosures about the brokerage firm and analyst research team and do not reflect analyst opinions. Following Li et al. (2024), I create a training set by sampling one report from each of the top 20 brokerage firms and manually labeling sentences. The fine-tuned BERT model achieves 97.31% accuracy on a held-out test set.
The Investext database provides detailed metadata for analyst reports, including analyst identities and release dates. Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and the matching window. I first match lead analyst names to unique analyst identifiers (AMASKCD) following the procedure in Li et al. (2024). I then associate each report with corresponding earnings forecasts, target prices, and stock recommendations in I/B/E/S. Since multiple reports may be issued between a forecast announcement and revision, I adopt the windowing strategy of Huang et al. (2014), linking reports to forecasts issued between two days prior to the I/B/E/S announcement date (ANNDATS) and two days after the revision date (REVDATS). [^4] I/B/E/S records both the forecast announcement date (ANNDATS) and the revision date (REVDATS), which marks when the forecast becomes outdated.
To isolate the distinct information content of analyst reports, I exclude reports released on earnings announcement days. The final matched sample consists of 122,252 analyst reports on S&P 100 stocks, authored by 1,305 analysts across 140 brokerage firms. Table 2 provides summary statistics of analyst reports across years and the Fama-French 12 industries, including the number of reports, unique brokerage firms, and analysts. The table also reports average report length, measured by page and token counts. The temporal distribution of reports follows the pattern documented by Bonini et al. (2023), with a peak in 2013 followed by a steady decline, likely driven by regulatory changes such as the Dodd-Frank Act and MiFID II (Bradshaw et al., 2017).
To quantify market reactions, I estimate stock-level price impacts using intraday data from the NYSE TAQ database for the out-of-sample period of January 2, 2015, to December 28, 2023. All monetary values are adjusted for inflation using the Consumer Price Index (CPI) and are normalized to December 2020.
Additionally, I match each analyst report with the latest earnings conference call transcript to control for concurrent events. Transcripts for S&P 100 firms are sourced from Seeking Alpha and matched to firm-quarter earnings announcement dates (RDQ) from Compustat, resulting in a sample of 7,909 firm-quarter earnings conference calls from 2005 to 2023.555Transcripts are available at https://seekingalpha.com/earnings/earnings-call-transcripts. A comparative analysis is presented in Section 3.2.2.
I include a set of firm-specific numerical features in the machine learning models, categorized as follows. Firm characteristics are sourced from CRSP and Compustat. [^6] Detailed variable definitions are provided in Table A2, and their summary statistics are in Table A3.
Report-Level Measures: $REC_{REV}$ denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm. $EF_{REV}$ represents earnings forecast revision, calculated as the current report’s EPS forecast minus the prior EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 trading days prior to the report’s release. $TP_{REV}$ measures target price revision, calculated as the difference between current and previous target prices in I/B/E/S for the same analyst-firm pair, scaled by the stock price 50 trading days prior to release. $Boldness$ is a binary indicator set to 1 when revisions are either above both the consensus and previous forecasts or below both. $SR$ denotes stock recommendations in the I/B/E/S rating system (1 = Strong Buy, 5 = Strong Sell). $ERet$ represents the 12-month return forecast, calculated as the 12-month price target scaled by the stock price 50 trading days prior to release. $PriorCAR$ measures cumulative abnormal return over [-10, -2] window preceding the report release.
Firm-Level Measures: $Size$ is quantified by market equity. Book-to-Market Ratio ($BtoM$) represents the ratio of book value to market value. Earnings surprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ denoting its absolute value. $Miss$ is a binary indicator equal to 1 if $SUE<0$. $TradingVolume$ represents the stock’s trading volume on recent earnings announcement days, standardized by shares outstanding. Distance to Default ($DD$) is sourced from the NUS Credit Research Initiative (CRI). $Fluidity$ measures firms’ product market competition following Hoberg et al. (2014).
Industry-Level Measures: $IndustryRecession$ is a binary indicator set to one if the stock’s Fama-French 48 industry experiences negative monthly returns in the bottom quintile.
3.2 Information Content of Analyst Reports